Summary of the GMVA Session, May 14-18, 2015 ============================================ The recording mode was at 2 Gbps using the PFB mode in dual polarisation. Effelsberg uses the RDBE recorder, Onsala and Pico Veleta record with the DBBC2. PdB records only half of the bandwidth, with its old MK5A system. Onsala participates with its new dual polarisation receiver. Yebes could not participate in this session, due to a broken 3mm receiver. The 3 KVN stations participated for some time in the 2nd part of the run (c151b). The GBT participated in some projects in part2 (c151b). The weather in the US was variable, with periods of overcast sky and heavy clouds at several stations. The following 16 stations observed in this run: Eb, On, Pv, PdB, GBT, 8xVLBA, 3xKVN General Summary: ================ Effelsberg and Onsala and the VLBA observed the full session, Pico Veleta and Plateau de Bure observed only some projects, as the IRAM TAC did not like all projects. The GBT observed in part 2 (c151b). Onsala has swapped polarisation. The weather at Pico Veleta was very good, at Effelsberg and Onsala it was good most of the time, and still acceptable at other times, the weather at PdB was marginal. For GBT and VLBA detailed information is not available, but see below. ----------------------------------------- Detailed notes: ftp fringe test: c151a: 134-1800 Eb-PV OJ287 SNR 14 t=7s c151b: 135-1300 Eb-On OJ287 SNR 24 t=5s, polarisation of ON swapped Effelsberg: recording on RDBE, c151a and c151b went well, mixture of good and bad weather May 14: 20 UT overcast but we can see and point on 3C273, Tsys 190K. May 15, 12 UT weather ok, Tsys 190K May 16, 08 UT overcast, but still can point on target 13-20 UT: very cloudy May 17, 00 UT cloudy, but pointing on 3C273/M87 ok 11 UT cloudy 19 UT better, ptg on OJ287 ok May 18 0 UT clear, thin haze, lost about 1hr on 0716+71 (ptg offset) 8 UT sun and clouds Pico: c151a and c151b went well, good weather all the time, no scans lost due to problems, except scans n0023 and n0029 on part b due to elev 85 deg constraint. Tsys ~ 140 K on May 14, 20 UT. Tsys ~ 125 K (8mm pvw) on May 15, 13 UT. PdB: reference antenna pad is N07 c151a: weather bad, only 1st 9 scans observers, problem with 1pps (10 MHz), ftp-fringe test negative Possible reason: bad cable connector on the 5 MHz distribution, repaired for c151b c151b: start on May 16, 23:45 UT due to bad weather before this time observed with 5 antennas phased, Tsys 90-110 K Onsala: c151a: DBBC core boards D and C not properly set, no fringes c151b: fringes found, polarisation is swapped, Tsys ~ 250 K in blue sky. We have recorded all the scans since the beginning of C151B. We had Pcal on and stable fmout-gps all the time. Before the 3mm session, our engineer told me the LO signal was locked to the right frequency while a bit of weak. This will be improved by replacing it with a new one in the future. After the experiment, I checked our 3mm receiver with the test signal. The LO signal looked OK. As far as the pointing, I used the pointing model suggested by our single-dish observation expert. From the operator log we see that C151B went smoothly although the weather was not always good for the 3mm observations. We had poor weather on Saturday, 16 May, raining and cloudy. Yebes: did not participate, dewar is broken KVN: observed for some time (US, YS, TN): c151b-a: May 16 01-05 UT no TN, Hexapod problem c151b-b: May 17 00-09 UT c151b-c: May 17/18 22-05 UT GBT: May 16, 20 UT: At Green Bank the sky was mostly clear last night when the run started. Tsys about 130K. There were some software problems at the startup, but it has for the most part run smoothly. Weather has been getting worse with a little rain just now. May 18: 20 UT: The recent 86 GHz VLBI run, C151B, aka GMVA14B_428, was completed at Green Bank. Most scans were recorded. The weather was not very good. Probalby the first few hours were ok, with some usable parts later now and then. Tsys started at about 133 K, but for some of the time was 160-200, and much of it had > 230 K. added Jun 3, 2015: the cal measurements for GBT are here https://safe.nrao.edu/wiki/bin/view/GB/Observing/WbandVLBACal/C151B VLBA: part1: VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C151A Observer: Krichbaum, T. Project type: OTHER Obs filename: c151acrd.* Date/Day: 2015MAY14/134 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is A. Sowinski. UT1-UTC -0.64; vclock okay. GPS delta ok except at %PT & MK where station timing disagrees with GPS by TIME roughly 1 second, but station timing checks ok via vclock. %KP Windy. 2057 Begin, except at MK. 2058 *NL Raining. 2140 New operator is Alan Kerr 2146 %OV windy. 2154-2156 *KP autostow. power glitch. repointed. COM 100 2 2357-0621 %NL precipitation 0000 Date is 2015MAY15/135. 0027 Begin at MK 0027 %MK raining. 0112 %FD raining. 0548 New operator is Betty Ragan. 0621 End. Downtime Summary: Total downtime : 2 min Percentage downtime of observing: 0.0% Average downtime per hour : 0.0 min Total scheduled observing time (# Antennas): 4512 min (8) part2: VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C151B Observer: Krichbaum, T. Project type: OTHER Obs filename: c151bcrd.* Date/Day: 2015MAY15/135 Ants Scheduled: NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is A. Sowinski. UT1-UTC -0.64; vclock okay. GPS delta ok except at %PT & MK where station timing disagrees with GPS by TIME roughly 1 second, but station timing checks ok via vclock. 1726 Begin, except at GB. 1726 PT, KP Windy. 1742 %NL Raining. 1759 %LA Raining. 1800 LA Windy. 2145 New operator is Alan Kerr 0000 Date is 2015MAY16/136. 0155 %PT raining. 0422 Begin at GB 0553 New operator is Betty Ragan. 0558-0559 *KP wind stow. WEA 100 1 0600-0602 *KP wind stow again. WEA 100 2 0607-0623 *KP stowed again. Leaving it out for a bit. WEA 100 16 0708-0709 *KP autostowed. Power glitch. COM 100 1 0709-0801 *PT, MK removed to run daily USNO observation. SCH 200 104 0709-0712 *KP wind stow. WEA 100 3 0719 %FD raining. 0801-0804 *MK, LA VMEs rebooted. Will need to re-run the SCH 100 3 USNO using these sites. VMEs need to be rebooted to get them to accept the new OBSERV.TX file. 0801-0959 *PT M401 is not running. Site tech is on his way HWC 100 118 out to reset it. (C136462) 0818-0820 *LA rebooted again. OBSERV.TX was not correct HUERR 100 2 last time. 0840-0842 *KP autostowed. Power glitch. COM 100 2 0844 %KP, MK raining. 0855-0947 *LA, MK removed for re-run of USNO observation SCH 200 104 that failed previously. 0900-2315 *KP 7mm cryo is warming up. (C136463) CRYO 50 427 0947 PT site tech has arrived. Parking antenna so he can work on the M401. 0959 PT M401 has been reset, and all looks well with it now. 1208-1217 *KP stowed for site tech to work on cryo. CRYO 100 9 1218 %PT raining. 1236-1246 *KP stowing antenna to check on the cryo again. CRYO 100 10 1252-1300 *KP stowed once more for cryo work. CRYO 100 8 1345 Operator is A. Sowinski. 1347-1436 *KP Site tech working on antenna again (C136463). CRYO 100 49 7mm cryo cooling now. 1600-1730 *KP A compressor failed. Now 3mm is warming as well. CRYO 50 45 Called site tech again (C136464). 2325-2326 *KP Autostow, power glitch. COM 100 1 2341-2342 *KP Autostow, power glitch. COM 100 1 0000 Date is 2015MAY17/137. 0140 New operator is Betty Ragan. 0140-0142 *GB DBE-1 had a timestamp sync error. Went TIME 100 2 away on its own. 0810 %BR raining. 0940-1009 *PT, MK removed to run daily USNO observation. SCH 200 58 0958-0959 *LA rebooted. Once again, we are going to have to SCH 100 1 run the USNO at LA instead. Rebooting the VME to set that up. 1001 %NL raining. 1009-1104 *LA removed for USNO re-observation. MK is also SCH 200 110 still being used for this version. 1009-1332 *PT is being put back into this observation, but HWC 100 203 the M401 is dead again. Will need to get the site tech back out to reset it, and it's going to take him a while. (C136466) 1104-1106 *LA had timing error. Had to resync and TIME 100 2 restart executor. 1346 Operator is A. Sowinski. 1928 OV windy. 0000 Date is 2015MAY18/138. 0137 New operator is Betty Ragan. 0602 %BR raining. 0805 End at GB. 0827 End. Downtime Summary: Total downtime : 1282 min Percentage downtime of observing: 4.2% Average downtime per hour : 2.5 min Total scheduled observing time (# Antennas): 30248 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.